The steam along the shelf break with a more open and thinner ice pack allowed the Palmer to move steadily towards the northern end of the survey grid. News from the Gould this evening is that they have been able to steam to the north at 3 to 5 kts for a good portion of the day and if their progress continues, they plan to work near the ice pack edge on the northern end of the survey grid on 24 August. This raises the possibility that they may not need our assistance to get to Palmer Station. Such a possibility has caused us to rethink our plans for the remainder of the cruise in a process that is still taking place. Our current position at 2203 on 23 August is -67° 05.077S; -73° 12.190W. Winds are out of the southwest (210) at about 17 kts and the air temperature is -18.5°C. The barometer is holding steady at 970.9 mlb.
The 22nd of August was a spectacularly bright day. Winds were light and there were few clouds most of the day. The visibility was really unlimited. Even though the sun does not get very high in the sky at this time of year, the light reflecting off the white ice pack, can make it very bright. Air temperatures remained cold, around -15°C. A trackline to take the Palmer to the L.M Gould's position was laid out to enable us to briefly sample at stations along the outer margin of the continental shelf along the southern portion of the survey grid. A science meeting was held at 1300, which was well attended, to discuss the scientific tasks that could be accomplished in the time available and where along the trackline they would be done. A flexible work plan was created that would allow for some work to be done over the next couple of days, while spending most of the time steaming.
We completed work at stations #82, 80, and 71, as well as finishing up the work at station 87 during the late night hours of 21/22 August. An XBT was dropped as we steamed through the stations #82 location. At station #80, only a CTD profile to 800 m and phytoplankton net tow were done. A more complete set of tasks was scheduled for station #71 and most, but not all were accomplished. These included a CTD, phytoplankton net tow, ice collection, 1-m2 and 10-m2 MOCNESS tows, a Tucker Trawl, and an ROV survey. The 1-m2 MOCNESS had electrical problems as described below and the 10-m2 MOCNESS tow was cancelled. Bird and mammal surveys were conducted for a large fraction of the daylight hours as a result of the increased amount of steaming time versus time spent working on station.
John Klinck reports on 22 August, the CTD group did one XBT and one CTD cast today. A second CTD cast bridged the change of day so its results will be reported tomorrow.
Station #82 (460 m) was measured en passant with an XBT. The first probe failed, but a second was launched which successfully made the full journey to the bottom. The cast showed a uniform mixed layer to 80 m and, below the thermocline, a deeper region of uniform temperature consistent with cold shelf water.
Station #80 (1850 m) is on the shelf break on the 140 line. The surface mixed layer has a slight salinity structure and extends to 140 m. The pycnocline extends over the next 80 m with only a few indications of layers. A broad temperature max extends from 280 to 420 m depth and shows some reversals of temperature. An oxygen minimum occurs between 290 and 350 m. Only the weakest of salinity maximum is observed in the deeper water properties.
Kendra Daly reports that in the last few days, the Daly group completed a Tucker trawl and two plummet net series, one in conjunction with acoustic data collection. During the Plummet net operation at station #83, there were sustained 45-50 knot winds with higher gusts, making it an exciting time for all. A few furciliae and one adult Euphausia superba, a few ctenophores, and several species of copepods were collected in the 0-100 m tow. The 100-320 m depth interval was dominated by Calanoid copepods.
At station #87, the Tucker trawl sampled to about 100 m depth and collected several ctenophores, a few E. superba furciliae, the euphausiid, Thysanoessa macrura, several kinds of copepods, pteropods, ostracods, and polychaete worms. The Hydroacoutics Technology Inc. (HTI) acoustic system was deployed at the same time as the Plummet net. Both the 120 and the 38 kHz frequencies indicated that a layer of krill was present between 90 and 100 m. The targeted Plummet net haul between the surface and 150 m confirmed that the layer was composed of adult krill. In addition, a few furciliae, Thysanoessa, ostracods, and some copepods were collected. The 150-250 m haul collected a few furciliae, some Thysanoessa, amphipods, ostracods, pieces of a siphonophore, small organge polychaetes, pteropods, and many copepods, such as Calanoides acutus, Calanus propinquus, Paraeuchaeta antartica, Metridia spp., and several small (ca. 1 mm) individuals that may be Undinella sp. or Scolecithricella. Some males and females of Calanus acutus and Calanoides propinquus are now present in samples. The deepest haul between 250 and 400 m had a similar composition to that of the intermediate depth haul.
The dive between station #75 and 76 revealed few larval krill on the broad open undersurface of the sea ice, but there were small aggregations up inside crevices. There were relatively few ctenophores, which were estimated to be about 1 every 10 m along a transect.
Egestion rates and assimilation efficiencies are being measured with the furciliae collected during the dive and the adults collected in the Plummet net. In addition, furciliae were collected from the 1-m MOCNESS net hauls at several depths to compare the size, dry weights, carbon, and lipid content of larvae in the water column with those collected under the ice. The results of growth and molting experiments indicate that these rates continue to remain relatively low.
Ari Friedlaender reported that as a result of the daylight hours on 21 August being spent either on station, or trawling nets around station #87, no proper marine mammal survey effort was made. However, incidental watch was kept from 0900-1200 as the Palmer steamed back and forth over a 2-mile area for towing purposes. Visibility was good, skies were cloudy, and the ice was 10/10 or ridged floes and numerous leads that were frozen over with new gray ice and grease ice. No cetacean sightings were made, but four crabeater seals were counted.
On 22 August, marine mammal observations were made from 0830-1530 as the Palmer steamed to station #82 and then 80. Visibility was excellent, with minke blows visible more than 3 nautical miles away. Skies were partly cloudy to mostly clear. Ice conditions were 10/10 for the most part. However, the Palmer steamed through a network of leads from 0930 on, which had pockets of open water mixed with leads of new gray ice and finger rafting nilas. Crabeater seals were seen in nearly all of the leads that we passed through. A total of 19 seals were counted in all. Several penguins were sighted as well throughout the transit (see bird observations). From 1345-1500, four cetacean sightings were made. All of the sightings occurred in the vicinity of a large pool-shaped lead, but none of the whales actually entered the open water. Blows were seen at varying distances from 1-3 nautical miles, but no bodies were seen for positive species identification. Due to the size and shape of the blow, as well as the habitat in which they were seen, the sightings were all logged as 'probable minke whale'. Two of the sightings were of single animals, while the other two were of at least two animals each, for a total of 6 whales seen. The blows were easily visible as they froze in the cold air and remained suspended for several seconds before being swept away in the wind. This was an encouraging sign because much of the open water was covered in patches of sea smoke that we thought might hinder our ability to discern a whale blow from simple rising fog. This was not the case, however, as the blows were far more substantial and easily recognized. Effort ended as we stopped at station 80 at 1530 (68° 17.36S, 75° 40.05W). A very good day for whale observations. It is remarkable that these whales are able to weave their way this deep into the pack ice, presumably following leads and bands of open water that are ever shifting in space and time. Are these whales moving into the ice from the ice edge and back, or are they meandering around completely in the ice? On the present course of action, there should be some steaming time around the ice edge to the north of the present location, and hopefully there will be some survey time around Palmer Station to see if there are whales in the open areas and thinner ice covered areas.
Ana Sirovic reported that on 22 August, she deployed a total of 3 difar sonobuys. The first one was deployed 5.5 nm from station #82 and was monitored for 1 h 49 min (over 10 mi range). The second one was deployed after the whale sightings and was monitored for 2 hours 55 min. The last buoy to be deployed failed immediately upon deployment. No biological noise was heard on any of the sonobuoys.
Chris Ribic and Erik Chapman report that on 21 August (JD233), there
was no surveying done because daylight was spent on station #87. There
was 7 hours of survey time on 22 August (JD-234) as the Palmer traveled
between stations #87, 82, and 80. The transit took the Palmer up
to the shelf break where there was more open water, with ice around 8/10ths
concentration. Over the shelf, the ice was at 9 and 10/10ths concentration.
Snow Petrels were the most abundant bird. Also seen were a small number
of Adélie and Emperor Penguins, Southern Giant Petrels, and Antarctic
Petrels. A summary of the observations is the following:
|Southern Giant Petrel||1|
BIOMAPER-II/MOCNESS report (P. Wiebe, C. Ashjian, and S. Gallager):
BIOMAPER-II was not deployed on 22 August.
A 1-m2 MOCNESS tow was conducted at station #71. The tow was fraught with technical difficulties. During preparation of the net for deployment, the ground wire was disconnected from the termination and required the installation of a new spade connector. During testing, it was found that the flow meter did not work. Two flow meters later, there was a functioning one. The net was deployed in over 1000 m of water at the edge of the continental shelf and slope. Unfortunately, the pressure sensor did not function when deployed and did not regain function when sent to depth where thawing usually occurs. The pressure reading from the OPC mounted on the MOCNESS had to be used to obtain depth. Because the OPC pressure sensor does not operate below 650 m or so, the net was deployed only to 650 m. The ice conditions were marginal and worsened during the tow, with multiple snags of the wire on ice floes in the wake of the ship occurring. Just as the wire snagged a persistent floe and required the ship to stop, communication with the MOCNESS was lost, but not the OPC. The message was that the "wire was open". After the wire was cleared of the ice floe, hauling the net up was continued while attempting to restart the MOCNESS data acquisition program. Wire snags on ice floes also continued during this period, and there were no data coming in from the MOCNESS and no way to trip any of the nets. Furthermore, the OPC data were not very useful because the OPC needs the flow meter data from the MOCNESS. The worsening towing conditions (ice snags) and the lack of meaningful information from the net, forced the cessation of the tow. It was brought on board as quickly and safely as possible. The ship was slowed to less than a knot and the net brought up at 30 m/min. It was recovered safely, with no damage. The loss of communication was caused by a broken ground wire. It is not yet known why the pressure sensor failed to work.
Two oblique tow samples were collected between 650 m and the surface.
Net 0 was preserved in ethanol for genetic analysis. Net 1 was preserved
in formalin for species information. Net 0 collected copepods, chaetognaths,
Thysanoessa, and euphuasiids. Net 1 collected chaetognaths and copepods
and 1 fish.